# Are Large Language Models Aligned with People’s Social Intuitions for Human–Robot Interactions?

### v1

@@ II BACKGROUND
In recent years, LLMs have started to become the focus of large parts of AI research. LLMs are pre-trained on massive text corpora scraped from the internet, using the causal language modeling objective in which the model predicts the next token given a left-sided context. Chat-variants like ChatGPT [10] or LLaMA-2Chat [11] are further fine-tuned to follow instructions by being trained on specific instruction-following datasets in a supervised manner, followed by being trained to better align with human preferences given multiple possible text completions, using the reinforcement learning from human feedback (RLHF) paradigm. LLMs also started playing a role in robotics as their encoded world knowledge allows them to suggest action plans without having to be fine-tuned on specific tasks, thus skipping much of the manual labor and domain expertise required in approaches like planning. Recent research uses LLMs not only to construct high-level plans that are then used to guide the robot’s behavior [3], but even to generate low-level motor commands given that control datasets were used for further training [4]. Given the current trend of integrating LLMs into robotics, these models are bound to play a role in coordinating a robot’s social behavior and interactions with humans [12, 13, 14]. This might happen through directly controlling the robot’s actions or by modeling users and their mental states to facilitate cooperation. Williams et al. [12] highlight the potential use of LLMs as placeholders in HRI-related robot architectures before more robust solutions can be developed. At the same time, they highlight the perils of using LLMs in HRI, referring to well-known problems, especially generating wrong statements (“hallucinating”), toxic text, or answers reflecting biases or stereotypes.
@@ IV-F MODELS
We use some of the most recent open- and closed-source models available. For open-source models, we choose the RLHF (chat) variants of LLaMA-2 with 13 billion and 70 billion parameters [10]. Not only is it easier to make RLHF models follow input instructions, but they outperform their base variants on various tasks. For closed-source models, we choose GPT-4 (API identifier gpt-4-0613), GPT-3.5 (gpt-3.5-turbo-0613), and the GPT-3 base model (davinci-002) [11, 32]. The GPT-3 base model is not trained with RLHF, contrasting the rest, all being trained to follow instructions. Lastly, Section V-D compares these text-only LLMs with the GPT-4 variant with vision capabilities (gpt-4-vision-preview, Feb. 2024), not relying on textual scenario descriptions but taking video frames as input.
@@ IV-G TECHNICAL DETAILS
To make the results reproducible, we use a greedy sampling approach, always choosing the most likely next token. This is achieved by setting the $temperature$ to 0 in the OpenAI API and the $top_{k}$ parameter to 1 with HuggingFace.
@@ V-D RESULTS: VISION–LANGUAGE MODELS
As an initial experiment, we test whether GPT-4 with vision correctly parses the videos and understands their content by generating descriptions of each situation. Simply understanding what happens in each video is a prerequisite for correctly assessing the value of different communicative acts. Manually analyzing the generated descriptions, we find that GPT-4 with vision correctly parses 50% of the videos. Among the correctly parsed situations are multiple ones primarily relying on dialogue as well videos relying on physical actions, for example, the robot knocking something over while trying to grasp it or the robot encountering an out-of-order sign. Moreover, the VLM also notes certain social cues, such as the user smiling in response to a successfully told joke. On the other hand, the VLM misinterprets multiple videos. Among others, when processing videos of suboptimal joint movements or pathfinding, the model simply notes the success at the end but does not mention the inefficiency of the solution. Furthermore, it fails to connect the dialogue to more static videos, claiming they are unrelated. Moreover, it misses the social norm violation of the robot forcing people to step out of its way when driving too closely to them and interprets the image of an empty box as a successful grasp instead of encountering a missing item.
@@ VI-A LIMITATIONS
Another limitation is that the LLMs were presented with only one scenario and one construct at a time when queried. Thus, models cannot rate scenarios or constructs relative to each other — a strategy a human participant is likely to adopt. However, when querying current models to answer a large amount questions at once, we observed worse answers, with the model often getting stuck in a loop of repeating the same answer. Future research should further analyze how different prompting techniques, such as the number of situations and constructs presented, influence model answers.
@@ VII-A1 TEXT-ONLY EXAMPLES
  • After being asked for a drink, the robot announces it is not sure where to find it.
@@ VII-A2 VIDEO EXAMPLES
In the VLM experiments (Section V-D), the model processes a sequence of video frames and a dialogue transcript. To illustrate the type of video, example frames from four videos are presented in Figure 6.
@@ VII-B EXPERIMENT 2 SCENARIO EXAMPLES [6, 7]
  • A robot waiter is serving alcoholic beverages at a party. When a man asks for another beer, it declines, saying “Sir, you have had enough.”

### v2

@@ II BACKGROUND
In recent years, LLMs have started to become the focus of large parts of AI research. LLMs are pre-trained on massive text corpora scraped from the internet, using the causal language modeling objective in which the model predicts the next token given a left-sided context. Chat-variants like ChatGPT [10] or LLaMA-2Chat [11] are further fine-tuned to follow instructions by being trained on specific instruction-following datasets in a supervised manner, followed by being trained to better align with human preferences given multiple possible text completions, using the reinforcement learning from human feedback (RLHF) paradigm. LLMs also started playing a role in robotics as their encoded world knowledge allows them to suggest action plans without having to be fine-tuned on specific tasks, thus skipping much of the manual labor and domain expertise required in approaches like planning. Recent research uses LLMs not only to construct high-level plans that are then used to guide the robot’s behavior [3], but even to generate low-level motor commands given that control datasets were used for further training [4]. Given the current trend of integrating LLMs into robotics, these models are bound to play a role in coordinating a robot’s social behavior and interactions with humans [12, 13, 14]. This might happen through directly controlling the robot’s actions or by modeling users and their mental states to facilitate cooperation. Williams et al. [12] highlight the potential use of LLMs as placeholders in HRI-related robot architectures before more robust solutions can be developed. At the same time, they highlight the perils of using LLMs in HRI, referring to well-known problems, especially generating wrong statements (hallucinating/confabulating), toxic text, or answers that reflect biases or stereotypes.
@@ IV-F MODELS
We use some of the most recent open- and closed-source models available. Firstly, we choose the RLHF (chat) variants of LLaMA-2 with 13 billion and 70 billion parameters [10]. Not only is it easier to make RLHF models follow input instructions, but they outperform their base variants on various tasks. Moreover, we choose GPT-4 (API identifier gpt-4-0613), GPT-3.5 (gpt-3.5-turbo-0613), and the GPT-3 base model (davinci-002) [11, 32]. The GPT-3 base model is not trained with RLHF, contrasting the rest, all having undergone instruction-finetuning and RLHF. Lastly, Section V-D compares these text-only LLMs with the GPT-4 variant with vision capabilities (gpt-4-vision-preview, Feb. 2024), not relying on textual scenario descriptions but taking video frames as input.
@@ IV-G TECHNICAL DETAILS
To make the results reproducible, we use a greedy sampling approach, always choosing the most likely next token. This is achieved by setting the $temperature$ to 0 in the OpenAI API and the $top_{k}$ parameter to 1 with HuggingFace. While the OpenAI API still suffers from some non-determinism, we verified that the variance in answers is minimal and does not affect the overall results.
@@ V-D RESULTS: VISION–LANGUAGE MODELS
As an initial experiment, we test whether GPT-4 with vision correctly parses the videos and understands their content by generating descriptions of each situation. Simply understanding what happens in each video is a prerequisite for correctly assessing the value of different communicative acts. Manually analyzing the generated descriptions, we find that GPT-4 with vision correctly parses 50% of the videos. Among the correctly parsed situations are multiple ones primarily relying on dialogue as well videos relying on physical actions, for example, the robot knocking something over while trying to grasp it or the robot encountering an out-of-order sign. Moreover, the VLM also notes certain social cues, such as the user smiling in response to a successfully told joke. On the other hand, the VLM misinterprets multiple videos. Among others, when processing videos of suboptimal joint movements or pathfinding, the model simply notes the success at the end but does not mention the inefficiency of the solution. Furthermore, it fails to connect the dialogue to more static videos, claiming they are unrelated. Moreover, it misses the social norm violation of the robot forcing people to step out of its way when driving too closely to them; and it interprets the image of an empty box as a successful grasp instead of encountering a missing item.
@@ VI-A LIMITATIONS
Another limitation is that the LLMs were presented with only one scenario and one construct at a time. Thus, models cannot rate scenarios or constructs relative to each other — a strategy a human participant is likely to adopt. However, when querying current models to answer a large amount questions at once, we observed worse answers, with the model often getting stuck in a loop of repeating the same answer. Future research should further analyze how different prompts influence model answers, e.g., through rephrasings or varying numbers of situations and constructs presented.
@@ VII-A2 VIDEO EXAMPLES
In the VLM experiments (Section V-D), the model processes a sequence of video frames and a dialogue transcript. Fig. 6 illustrates the type of videos.